OpenStudy (anonymous):

y varies inversely with x. Find the inverse variation equation if x = 12 when y = -18. a. y/x = -3/2 b. y/x = -216 c. xy = -216 d. xy = -3/2

OpenStudy (anonymous):

I dont gett itt

OpenStudy (anonymous):

@Abhisar

OpenStudy (anonymous):

i think i remember how these work inverse variation means \[y=\frac{k}{x}\] for some \(k\) (they always use k) and you have to find it

OpenStudy (anonymous):

\[y= -\frac{ 3x }{ 2 }\] ?

OpenStudy (anonymous):

would it be x/y?

OpenStudy (anonymous):

y/x *

OpenStudy (anonymous):

sorry site crashed for me \[y=\frac{k}{x}\] \[-18=\frac{k}{12}\]

OpenStudy (anonymous):

what does that mean? the -18= k/12

OpenStudy (anonymous):

it means that \[xy=k\] so we solve for \(k\) and get \(k=-18\times 12\)

OpenStudy (anonymous):

ohh ok